Compound amount


find the compound amount and the interest on 100,000 at 6% compounded semi-annually for

(a) 10 years

(b) 20 years

(c) 30 years


1
Expert's answer
2021-03-18T13:39:28-0400

(a) "A=100000(1+\frac{0.06}{2})^{10*2}=180611.12."

"I=180611.12-100000=80611.12."


(b) "A=100000(1+\frac{0.06}{2})^{20*2}=326203.78."

"I=326203.78-100000=226203.78."


(c) "A=100000(1+\frac{0.06}{2})^{30*2}=589160.31."

"I=589160.31-100000=489160.31."